OpenStudy (anonymous):

The number of solutions of the equation 3p + 4q = 70, where p and q are positive integers and p > q, is

OpenStudy (anonymous):

\[3p = 70 - 4q\] \[p = \frac{70 - 4q}{3}\] As p and q are positive integers so, Put q = 1, 2, 3 and so on so that the corresponding value of p is an positive integer.. Firstly put q = 1, \[p = \frac{66}{3} = 22\] Like wise you will get three values for p and q for which the equation holds true according to the condition, p>q. These values are: \[(p,q) = (22, 1), (4, 18), (7, 14)\] So, number of solutions will be : 3..

ganeshie8 (ganeshie8):

correcting small typo ;) These values are: (p,q)=(22,1),(18, 4),(14, 7)

OpenStudy (anonymous):

Oh really thanks...@ganeshie8 ..